OpenStudy (anonymous):

Find the roots of the polynomial equation. x2 + 24 = –11x

OpenStudy (linn99123):

x² + 24 = - 11x x² + 11/2x = - 24 + (11/2)² x² + 11/2x = - 96/4 + 121/4 (x + 11/2)² = 25/4 x + 11/2 = 5/2 x = 5/2 - 11/2, x = - 6/2, x = - 3 x = - 5/2 - 11/2, x = - 16/2, x = - 8

OpenStudy (anonymous):

so -8,-3 is the answer